OpenStudy (lilshane):

solve. 4/6 = 24/a a=

OpenStudy (lilshane):

@iGreen

OpenStudy (anonymous):

hmm

OpenStudy (lilshane):

@Frostbite

OpenStudy (lilshane):

@texaschic101

OpenStudy (texaschic101):

4/6 = 24/a --- this is a proportion, you can cross multiply (4)(a) = (24)(6) 4a = 144 a = 36 you will notice that proportions are nothing but equivalent fractions and 4/6 when reduced = 2/3 and when 24/36 reduce, you also get 2/3.....these are nothing but equivalent fractions

OpenStudy (lilshane):

thank

OpenStudy (texaschic101):

anytime :)